OpenStudy (anonymous):

Help plz 1) Find the equation of the line in point-slope form containing the point (-4,7) and parallel to the line 3x-5y=9

OpenStudy (saifoo.khan):

hey! its easy!

OpenStudy (anonymous):

1) present given equation in form : y=... 5y=3x-9 y=3/5 *x - 9/5 2) you found that the slope of given line is 3/5 your line = parallel to the give line, so - the same slope; 3) equation of the line containing point (-4,7): y-7=(3/5) *(x- (-4)) y=(3/5)x +(12/5) +7 y=(3/5)x + 47/5

OpenStudy (saifoo.khan):

YeaH! Correct. i was just about to type it. :)

OpenStudy (anonymous):

is this all its asking? cuz i got that answer too but i wasnt sure if it was finished

OpenStudy (saifoo.khan):

it fine. If i had to write it thn, i would hav written: 5y=3x +47.

OpenStudy (anonymous):

Ok, but what if its a perpendicular line instead of a parallel line?

OpenStudy (anonymous):

use the reciprocal correct? which would be instead of 3/5 => -5/3?

OpenStudy (saifoo.khan):

thn v'll apply the formuale: m1 x m2 = -1 Yeah! crrect,

OpenStudy (anonymous):

so answer would be y = (-5/3)x +(1/3)

OpenStudy (saifoo.khan):

there should a cordinate through which u may find wht is the cordinate?

OpenStudy (anonymous):

um you mean the points its containing? its "...containing the point (-4,7) and perpendicular to the line 3x-5y=9"

OpenStudy (saifoo.khan):

oh! thn the gradient of this line will b 1/3.

OpenStudy (saifoo.khan):

thn u may add the value of y-intercept which u may find with a cordinate & the gradient 1/3

OpenStudy (anonymous):

i dont understand, what do u mean?

OpenStudy (saifoo.khan):

till where u understood?

OpenStudy (anonymous):

i got the answer of \[y = (-5 \ 3)x + (1 \ 3)\] <

OpenStudy (anonymous):

weird... those two are fractions* -5/3 and 1/3

OpenStudy (saifoo.khan):

this is the main line on which there is a perpendicular?

OpenStudy (anonymous):

i believe so, thats what i solved for using the point slope formula. and i got the slope from the problem i had done above with the parallel line

OpenStudy (saifoo.khan):

sorry man! i gotta go now. tc. :)